OpenStudy (anonymous):

for the arithmetic sequence determine the t1 and d : _,4,_,_,10 ,the answers 2,4,6,8,10 but how do you get t1 given only t2=4,t5=10?

OpenStudy (amistre64):

we know it arithmatic right?

OpenStudy (amistre64):

so it has to follow a formula for arithmatic seqs

OpenStudy (amistre64):

count 4 as your starter and use sequence it as this to find d an = a1 + d(n-1) a4 = 4 + d(4-1) 10 = 4 + d(3) 6/3 = d = 2 an = a1 + 2(n-1)

OpenStudy (amistre64):

the difference is 2 apart; 4 - 2 = 2; 4+2 = 6, 10-2 = 8

OpenStudy (anonymous):

THANKS SO MUCH =D

OpenStudy (amistre64):

i hope the process did you good :)

OpenStudy (amistre64):

theres prolly a more cogent way to go about it tho ...

OpenStudy (anonymous):

simple is refreshing :)

OpenStudy (amistre64):

:) good

OpenStudy (anonymous):

\[t _{n}=d \times n +c\] for t(2): d(2)+c=4 for t(5): d(5)+c=10 subtracting the first from the second gives 3d=6; d=2; c=0 now we can find any t(n)=2n